Find the greatest length of an organ pipe open at both ends that will have its fundament frequency in the normal hearing range (20-20,000 Hz). Speed of sound in air = 340 m s-1.

Grade:10

1 Answers

Jitender Pal
9 years ago
The resonance column apparatus is equivalent to a closed organ pipe. Here l = 80 cm = 10 × 10^–2 m ; v = 320 m/s ⇒n base 0 = V/4l = 320/4 * 50 * 10^-2 = 100 Hz So the frequency of the other harmonics are odd multiple of n base 0 = (2n + 1) 100 Hz According to the question, the harmonic should be between 20 Hz and 2 KHz.